# Boosting Adverse Weather Crowd Counting via Multi-queue Contrastive Learning

### v1

@@ IV-B NWPU-Weather Dataset
IV-C JHU-Crowd++ Dataset
As shown in Table III, despite the challenges posed by the diverse scenes, complex and variable weather conditions and weather class imbalance in the JHU-Crowd++ dataset, MQCL achieves an improvement of 8.6% in MAE and 6.9% in RMSE under adverse weather conditions compared to the baseline ConvNeXt. This indicates that MQCL not only performs well on synthetic datasets but also effectively boosts the model’s robustness in real-world datasets. However, while MQCL achieved significant improvements over the baseline model ConvNeXt, its performance still slightly lags behind the current state-of-the-art algorithms. We summarize the reasons as follows: 1) as shown in Table III, all methods outperforming MQCL, including CLTR, MAN, and AWCC-Net, have introduced Transformer-based modules, the computational complexity of which is significantly higher than MQCL. The MAE of AWCC-Net under adverse weather conditions is 9.5% lower than our method, but it incurs over 85.9% more FLOPs; 2) MQCL focuses on the design of learning strategies. It only introduces a lightweight refiner module without significantly altering the architecture of the backbone network. The visualizations of ConvNeXt and our MQCL are compared in Fig. 6.
TABLE V: Quantitative results comparing with the state-of-the-art methods on the GCC dataset. The numbers in parentheses represent the rankings of our method.
@@ IV-C JHU-Crowd++ Dataset
IV-D GCC Dataset
GTA5 Crowd Counting (GCC) dataset [30] is a fully synthetic large scale dataset, the scenarios of which is derived from the computer game Grand Theft Auto V (GTA5). The dataset was initially proposed for pretraining the model and boosting the performance on the real-world datasets by finetuning or adaptation. Here, we use it as a benchmark to validate the performance of crowd counting models under adverse weather conditions on large-scale data.
There are 15212 images and 7.63 million labels contained in the GCC dataset. Since each location in the dataset encompasses multiple perspectives, [30] provides three strategies for splitting the training and the testing set: random, cross-camera and cross-location. Here we choose the cross-location strategy, which means 75%/25% locations are utilized for training/testing, since this scheme is most effective in validating the robustness of the model against unknown scenes. The weather conditions in the GCC dataset are diverse, comprising a total of seven types: normal, cloudy, rainy, foggy, thunder, overcast and extra sunny. They are evidently unevenly distributed either and the pie chart of the weather distribution is shown in Fig. 5 (b). In our implementation, we randomly selected 10 locations out of the 75 in the training set as a validation set. And it is worth noting that we choose ConvNeXt-S rather than ConvNeXt-T as our backbone model due to the large-scale nature of the GCC dataset.
The quantitative results of our method and several state-of-the-art methods are compared on the GCC dataset, as shown in Table V. As can be seen, our MQCL achieves optimal performance under both normal and adverse weather conditions on all the metrics. Comparing to our baseline ConvNeXt, MQCL decreases the MAE/RMSE under normal and adverse weather conditions by 8.6%/4.9% and 6.8%/5.1%, respectively. This indicates that our MQCL can also outperforms other methods when trained on relatively large-scale datasets and more importantly, our multi-queue MoCo strategy performs well even in the case of higher number of sub queues (7 for example).
IV-E Key Issues and Discussions
The significance of contrastive learning. As discussed earlier, the prerequisite for addressing the multi-domain learning problem is that the model can perceive different domains. If the encoder fails to extract distinctive weather information, the refiner will lack the target for conversion and struggle to fulfill its intended purpose. As shown in Table VI, if contrastive learning is not employed in the first stage to enable the model to learn weather information with discriminative characteristics, and the refiner module is directly introduced afterwards, the performance of the model actually decreases significantly under both normal and adverse weather conditions. We believe this is because the refiner module, trained from scratch without clear targets, may disrupt the crowd information extracted by the original encoder.
TABLE VI: The significance of contrastive learning in our method. “Refine w/o CL” adopts a similar approach to MQCL, except that $\mathcal{L}_{contra}$ and $\mathcal{L}_{supcon}$ are removed from the loss functions, respectively.
@@ IV-E Key Issues and Discussions
TABLE VII: Performance comparison of the two optional strategies on the JHU-Crowd++ dataset. Strategy 1 treats samples with the same weather label as positive examples and strategy 2 treats samples originating from the same image as positive examples.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
Strategy 1 51.4 153.0 99.7 540.4
Strategy 2 (Ours) 50.5 152.0 96.5 522.8
Strategy of positive/negative selection. As mentioned above, there are two optional strategies in the WRL stage: 1) treating samples with the same weather label as positive examples; 2) treating samples originating from the same image as positive examples; We conduct experiments on the JHU-Crowd++ dataset to compare these two strategies. The model performance after the WRL stage using these two strategies is shown in Table VII. Strategy 2 outperforms Strategy 1 in both normal and adverse conditions. We attribute this to the following reasons: 1) even if two samples share the same weather label, their weather conditions may still vary significantly. The practice of minimizing all the representations with the same label is not in line with the target of us; 2) samples from the same image not only share the same weather conditions but also the same scene. Strategy 2 has the potential to enhance the model’s scene recognition capabilities.
@@
TABLE VIII: Performance comparison of the three storage strategies for vectors K on the JHU-Crowd++ dataset.
@@
The storage strategy for vectors K. There are three optional storage strategies for vectors K: 1) memory bank [34]; 2) single-queue MoCo [7]; 3) multi-queue MoCo (ours). Extensive experiments are conducted on the JHU-Crowd++ dataset to compare the above strategies. The t-SNE [35] visualization of the vectors Q after the WRL stage on the JHU-Crowd++ dataset is shown in Fig. 7. The memory bank strategy suffers from a severe lack of discriminative capacity. The representations learned by the single-queue strategy are also not discriminative enough, especially near the rain weather representations in Fig. 7(b). The multi-queue strategy does not suffer from the aforementioned issues. The significantly dispersed clusters shown Fig. 7(c) illustrate that the properties learned from the multi-queue MoCo represent weather information well. The performance comparison shown in Table VIII also demonstrates that the proposed multi-queue MoCo can effectively address the class imbalance problem. We attribute the phenomenon to the following reasons: 1) while the memory bank can store a large number of samples with minimal memory consumption, it does not employ a stable strategy to update the encoder, and the sample update frequency is too low, resulting in poor sample consistency; 2) although single-queue MoCo addresses the issue of poor sample consistency by introducing a queue and momentum update strategy, the class imbalance problem in the dataset leads to inconsistencies between the data distribution in the queue and the uniform distribution prior; 3) multi-queue MoCo not only retains the advantages of high sample consistency but also greatly alleviates the problem of inconsistency between data distribution and the uniform prior. Thus, it achieves the best performance.
TABLE IX: Ablation study on the JHU-Crowd++ dataset.
@@
Ablation study. Ablation studies are performed on the JHU-Crowd++ dataset and the quantitative results are shown in Table IX. We start with the baseline of the end-to-end model, i.e., only the encoder and decoder. First, the effectiveness of multi-queue MoCo is tested. An improvement of 2.8%/1.5% and 7.5%/6.6% in MAE/RMSE under normal and adverse weather is achieved compared to the baseline. From this, it can be seen that most of the performance improvements under adverse weather conditions come from the representation learning in the WRL stage. This verifies the effectiveness of the proposed multi-queue MoCo for enhancing robustness under adverse weather conditions. Additionally, the performance improvement under normal weather conditions corroborates the earlier analysis that the strategy that treats different augmentations from the same image as positive samples can aid in strengthening the scene recognition capabilities of the model. Subsequently, the refiner is added, and the best performance is achieved, with a reduction of 1.4%/0.4% in MAE/RMSE under normal weather conditions and 0.7%/0.3% under adverse weather conditions, respectively. This demonstrates that, under the guidance of supervised contrastive learning, the refiner is capable of converting adverse weather representations to the normal domain, enabling the decoder to focus on a single domain, resulting in performance improvements under both normal and adverse weather conditions.
TABLE X: Comparison between multi-queue MoCo and other class imbalance contrastive learning strategies on the JHU-Crowd++ dataset.
@@
Comparison with other class imbalance strategies. In this section, we will discuss the strengths and weaknesses of multi-queue MoCo compared to the five representative imbalanced contrastive learning strategies mentioned in Section II-C, and investigate their applicability to our MQCL method. Attributed to MAK’s [40] strategy of sampling from the real world, the diversity of tail-end classes is significantly increased, the issue of class imbalance can be effectively addressed. However, it is not applicable to our crowd counting task because only unlabeled samples can be supplemented, but in crowd counting, head position labels are required for learning. The rest four methods can be applied to our model, we conduct experiments on the JHU-Crowd++ dataset to compare and analyze their performance results.
In the experiment setup, only $\mathcal{L}_{contra1}$ in the MQCL is replaced with the loss functions of the four imbalanced contrastive learning strategies, and all other settings remains unchanged. The results are shown in Table X. The performance of KCL [39] is worse than multi-queue MoCo because KCL samples a sufficient number of positive samples, but the negative samples remain imbalanced. SBCL [42] fundamentally alleviates the issue of class imbalance through the method of subclass partitioning. The experiments in [42] show that SBCL outperforms KCL [39] and TSC [43] in image classification tasks. However, it achieved the lowest performance in this experiment. We think the approach of dividing the normal weather class into multiple subclasses may bring confusion to the CRR stage. Because in the CRR stage, the representations are pulled towards the normal weather domain, which, however, has multiple clusters. BCL [41] combines gradient balancing and batch data balancing; TSC [43] introduces cluster center priors. They achieved performance almost as excellent as multi-queue MoCo. Although these three methods exhibit similar performance, they address the imbalance problem from different perspectives, each with its own significance.
TABLE XI: Generalization ability of ConvNeXt-T trained on the NWPU-Normal dataset and the NWPU-Weather dataset on the test set of the JHU-Crowd++ dataset.
@@
The significance of the NWPU-Weather dataset. As described in Section IV-B, we synthesize the NWPU-Weather dataset to validate the models’ performance under adverse weather conditions. However, beyond this, in terms of the significance of proposing this dataset, there is a question: how is the generalization ability of the model trained on the synthetic NWPU-Weather dataset? If the model trained on the NWPU-Weather dataset can’t be generalized to the real world, the contribution of the synthetic dataset is limited. Thus we conduct experiments to validate this question.
@@
As is shown in Fig. XI, compared to the model trained on the NWPU-Normal dataset, that trained on our synthesised NWPU-Weather dataset is 3.0%/11.3% and 1.3%/29.5% stronger in MAE/RMSE under normal and adverse weather, respectively. This validates the significance of our NWPU-Weather dataset: more diversified weather conditions in the training data imparts better robustness to the model. Especially notable is the improvement in RMSE under adverse weather conditions, indicating that the synthesized data particularly excels in enhancing the stability of the model.
V Conclusion and Limitations
In this paper, we propose a contrastive learning-based method called MQCL to tackle the problem of class-imbalanced adverse weather crowd counting and synthesize a new adverse weather crowd counting dataset. To address the dual challenges of image degradation and class imbalance, the multi-queue MoCo is employed to enable the model to learn weather-aware representations. Furthermore, a contrastive-learning-based refining method is utilized to guide the refiner on representation conversion. Extensive experiments are conducted to compare and choose the strategy of contrastive learning. We believe that this is not only applicable to the crowd counting task but also holds significant reference value for other domains.
MQCL has achieved significant performance improvements compared to the baseline. However, there are still some limitations. For example, we focus on the learning strategy and maintaining the model’s lightweight nature, but its performance still falls slightly behind the state-of-the-art models based on Transformers. Secondly, for this study, our focus is on the application of multi-queue MoCo in the adverse weather crowd counting task, but it is a generic method to deal with data imbalance, its performance in other tasks awaits future research. Additionally, we only utilized synthetic methods to construct the dataset, leading to a certain domain gap between real-world scenes.

### v2

@@ 4.3 JHU-Crowd++ Dataset
Table 3 shows the performance of MQCL and other state-of-the-art methods on the JHU-Crowd++ dataset. Despite the challenges posed by the diverse scenes, complex and variable weather conditions and weather class imbalance in the JHU-Crowd++ dataset, MQCL achieves an improvement of 8.6% in MAE and 6.9% in RMSE under adverse weather conditions compared to the baseline ConvNeXt. This indicates that MQCL not only performs well under synthesized weather conditions but also effectively boosts the model’s robustness in real-world weather conditions. However, while MQCL achieved significant improvements over the baseline model ConvNeXt, its performance still slightly lags behind the current state-of-the-art algorithms. We summarize the reasons as follows: 1) as shown in Table 4, all methods outperforming MQCL, including CLTR, MAN, and AWCC-Net, have introduced Transformer-based modules, the computational complexity of which is significantly higher than MQCL. The MAE of AWCC-Net under adverse weather conditions is 9.5% lower than our method, but it incurs over 85.9% more FLOPs; 2) MQCL focuses on the design of learning strategies. It only introduces a lightweight refiner module without significantly altering the architecture of the backbone network. The visualizations of ConvNeXt and our MQCL are compared in Fig. 6.
4.4 GCC Dataset
GTA5 Crowd Counting (GCC) dataset [43] is a fully synthetic large scale dataset, the scenarios of which is derived from the computer game Grand Theft Auto V (GTA5). The dataset was initially proposed for pretraining the model and boosting the performance on the real-world datasets by finetuning or adaptation. Here, we use it as a benchmark to validate the performance of crowd counting models under adverse weather conditions on large-scale data. There are 15212 images and 7.63 million labels contained in the GCC dataset. Since each location in the dataset encompasses multiple perspectives, [43] provides three strategies for splitting the training and the testing set: random, cross-camera and cross-location. Here we choose the cross-location strategy, which means 75%/25% locations are utilized for training/testing, since this scheme is most effective in validating the robustness of the model against unknown scenes. The weather conditions in the GCC dataset are diverse, comprising a total of seven types: normal, cloudy, rainy, foggy, thunder, overcast and extra sunny. They are evidently unevenly distributed either and the pie chart of the weather distribution is shown in Fig. 5 (b). In our implementation, we randomly selected 10 locations out of the 75 in the training set as a validation set. And it is worth noting that we choose ConvNeXt-S rather than ConvNeXt-T as our backbone model due to the large-scale nature of the GCC dataset.
Table 5: Quantitative results comparing with the state-of-the-art methods on the GCC dataset. The numbers in parentheses represent the rankings of our method.
@@ 4.4 GCC Dataset
The quantitative results of our method and several state-of-the-art methods are compared on the GCC dataset, as shown in Table 5. As can be seen, our MQCL achieves optimal performance under both normal and adverse weather conditions on all the metrics. Comparing to our baseline ConvNeXt, MQCL decreases the MAE/RMSE under normal and adverse weather conditions by 8.6%/4.9% and 6.8%/5.1%, respectively. This indicates that our MQCL can also outperforms other methods when trained on relatively large-scale datasets and more importantly, our multi-queue MoCo strategy performs well even in the case of higher number of sub queues (7 for example).
Table 6: The significance of contrastive learning in our method. “Refine w/o CL” adopts a similar approach to MQCL, except that $\mathcal{L}_{contra1}$ and $\mathcal{L}_{contra2}$ are removed from the loss functions, respectively.
@@
4.5 Key Issues and Discussions
Strategy of positive/negative selection. As mentioned above, there are two optional strategies in the WRL stage: 1) treating samples with the same weather label as positive examples; 2) treating samples originating from the same image as positive examples; We conduct experiments on the JHU-Crowd++ dataset to compare these two strategies. The model performance after the WRL stage using these two strategies is shown in Table LABEL:tbl:strategy_comparison. Strategy 2 outperforms Strategy 1 in both normal and adverse conditions. We attribute this to the following reasons: 1) even if two samples share the same weather label, their weather conditions may still vary significantly. The practice of minimizing all the representations with the same label is not in line with the target of us; 2) samples from the same image not only share the same weather conditions but also the same scene. Strategy 2 has the potential to enhance the model’s scene recognition capabilities.
@@ 4.5 Key Issues and Discussions
Table 7: Performance comparison of the three storage strategies for vectors K on the JHU-Crowd++ dataset.
@@
The storage strategy for vectors K. There are three optional storage strategies for vectors K: 1) memory bank [44]; 2) single-queue MoCo [8]; 3) multi-queue MoCo (ours). Extensive experiments are conducted on the JHU-Crowd++ dataset to compare the above strategies. The t-SNE [37] visualization of the vectors Q after the WRL stage on the JHU-Crowd++ dataset is shown in Fig. 7. The memory bank strategy suffers from a severe lack of discriminative capacity. The representations learned by the single-queue strategy are also not discriminative enough, especially near the rain weather representations in Fig. 7(b). The multi-queue strategy does not suffer from the aforementioned issues. The significantly dispersed clusters shown Fig. 7(c) illustrate that the properties learned from the multi-queue MoCo represent weather information well. The performance comparison shown in Table 7 also demonstrates that the proposed multi-queue MoCo can effectively address the class imbalance problem. We attribute the phenomenon to the following reasons: 1) while the memory bank can store a large number of samples with minimal memory consumption, it does not employ a stable strategy to update the encoder, and the sample update frequency is too low, resulting in poor sample consistency; 2) although single-queue MoCo addresses the issue of poor sample consistency by introducing a queue and momentum update strategy, the class imbalance problem in the dataset leads to inconsistencies between the data distribution in the queue and the uniform distribution prior; 3) multi-queue MoCo not only retains the advantages of high sample consistency but also greatly alleviates the problem of inconsistency between data distribution and the uniform prior. Thus, it achieves the best performance.
Table 8: Ablation study on the JHU-Crowd++ dataset.
@@
Ablation study. Ablation studies are performed on the JHU-Crowd++ dataset and the quantitative results are shown in Table 8. We start with the baseline of the end-to-end model, i.e., only the encoder and decoder. First, the effectiveness of multi-queue MoCo is tested. An improvement of 2.8%/1.5% and 7.5%/6.6% in MAE/RMSE under normal and adverse weather is achieved compared to the baseline. From this, it can be seen that most of the performance improvements under adverse weather conditions come from the representation learning in the WRL stage. This verifies the effectiveness of the proposed multi-queue MoCo for enhancing robustness under adverse weather conditions. Additionally, the performance improvement under normal weather conditions corroborates the earlier analysis that the strategy that treats different augmentations from the same image as positive samples can aid in strengthening the scene recognition capabilities of the model. Subsequently, the refiner is added, and the best performance is achieved, with a reduction of 1.4%/0.4% in MAE/RMSE under normal weather conditions and 0.7%/0.3% under adverse weather conditions, respectively. This demonstrates that the refiner is capable of converting adverse weather representations to the normal domain, enabling the decoder to focus on a single domain.
Table 9: Comparison between multi-queue MoCo and other class imbalance contrastive learning strategies on the JHU-Crowd++ dataset.
@@
Comparison with other class imbalance strategies. In this section, we will discuss the strengths and weaknesses of multi-queue MoCo compared to the five representative imbalanced contrastive learning strategies, and investigate their applicability to our MQCL method. Attributed to MAK’s [12] strategy of sampling from the real world, the diversity of tail-end classes is significantly increased, the issue of class imbalance can be effectively addressed. However, it is not applicable to our crowd counting task because only unlabeled samples can be supplemented, but in crowd counting, head position labels are required for learning. The rest four methods can be applied to our method, we conduct experiments on the JHU-Crowd++ dataset to compare and analyze their performance results.
In the experiment setup, only $\mathcal{L}_{contra1}$ in the MQCL is replaced with the loss functions of the four imbalanced contrastive learning strategies, and all other settings remains unchanged. The results are shown in Table 9. The performance of KCL [13] is worse than multi-queue MoCo because KCL samples a sufficient number of positive samples, but the negative samples remain imbalanced. SBCL [9] fundamentally alleviates the issue of class imbalance through the method of subclass partitioning. The experiments in [9] show that SBCL outperforms KCL [13] and TSC [18] in image classification tasks. However, it achieved the lowest performance in this experiment. We think the approach of dividing the normal weather class into multiple subclasses may bring confusion to the CRR stage. Because in the CRR stage, the representations are pulled towards the normal weather domain, which, however, has multiple clusters. BCL [46] combines gradient balancing and batch data balancing; TSC [18] introduces cluster center priors. They achieved performance almost as excellent as multi-queue MoCo. Although these three methods exhibit similar performance, they address the imbalance problem from different perspectives, each with its own significance.
Table 10: The performance of MQCL on the JHU-Crowd++ dataset under different encoder configurations.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
VGG [33] 68.2 221.4 115.3 632.0
VGG + MQCL 67.2 209.2 101.4 533.4
CSRNet [19] 64.3 223.5 114.8 603.9
CSRNet + MQCL 63.1 207.7 99.8 540.2
Swin [25] 57.8 171.8 103.5 571.7
Swin + MQCL 54.6 163.8 99.5 558.8
ConvNeXt [26] 52.7 154.9 105.1 561.4
MQCL 50.5 152.0 96.5 522.8
Performance under different encoder configurations. Since the focus of MQCL is the designing of learning strategies, it is necessary to validate the effectiveness of MQCL under different network architectures. We conduct experiments on the JHU-Crowd++ dataset, where the encoder of MQCL is replaced with other models, including simple CNN models like VGG [33] and CSRNet [19] and a Transformer-based model Swin-T [25]. The result is shown in Table 10. The structure of the base models (ConvNeXt, CSRNet, VGG, Swin) keep consistent of the original model, only the classification head is replaced with the same decoder used in MQCL. Clearly, applying any encoder structure to MQCL results in significantly better performance compared to the corresponding baseline. This demonstrates the versatility of the learning methods in MQCL.
Table 11: Generalization ability of ConvNeXt-T trained on the NWPU-Normal dataset and the NWPU-Weather dataset on the test set of the JHU-Crowd++ dataset.
@@
The significance of the NWPU-Weather dataset. We think the significance of the NWPU-Weather dataset is two-fold. Firstly, since there is currently only one fully real-world adverse weather crowd counting dataset, and the GCC dataset is not based on real-world scenarios. We think that these two datasets are insufficient to validate the model’s performance under adverse weather conditions. Therefore, we synthesize the NWPU-Weather dataset based on real-world scenarios as a supplement. Additionally, beyond this, there is a question: how is the generalization ability of the model trained on the synthetic NWPU-Weather dataset? If the model trained on the NWPU-Weather dataset can’t be generalized to the real world, the contribution of the synthetic dataset is limited. Thus we conduct experiments to validate this question.
@@
As is shown in Fig. 11, compared to the model trained on the NWPU-Normal dataset, that trained on our synthesised NWPU-Weather dataset is 3.0%/11.3% and 1.3%/29.5% stronger in MAE/RMSE under normal and adverse weather, respectively. This validates the significance of our NWPU-Weather dataset: more diversified weather conditions in the training data imparts better robustness to the model.
5 Conclusion and Limitations
In this paper, we propose a contrastive learning-based method called MQCL to tackle the problem of class-imbalanced adverse weather crowd counting and synthesize a new adverse weather crowd counting dataset. The core contributions of this paper are multi-queue MoCo and the contrastive-learning-based refining method, which respectively achieve weather-aware representation learning and representation conversion, addressing the dual challenges of image degradation and class imbalance. Our work has achieved significant performance improvements compared to the baseline but there are still some limitations. For example, we only utilized synthetic methods to construct the dataset, leading to a certain domain gap between real-world weather. Additionally, for this study, our focus is on the application of multi-queue MoCo in the adverse weather crowd counting task. But it is a generic method to deal with data imbalance and is not only applicable to the crowd counting task but also holds significant reference value for other domains. Its performance in other tasks awaits future research.
6 Acknowledgement
This work was partially supported by the National Natural Science Foundation of China (62476130, 62176123), and Research Fund of Guangxi Key Lab of Multi-source Information Mining & Security, China (MIMS23-M-03).